OpenStudy (anonymous):

Which identity could be the missing step to varify the identity (cos theta) (csc theta) = cot theta? a. Cos theta (1/sec theta) = cot theta b. cos theta (1/sin theta) = cot theta c. sin theta (1/ csc theta) = cot theta d. cos theta (1/ tan theta) = cot theta

OpenStudy (math&ing001):

The one that makes sense is cos theta (1/sin theta) = cot theta since csc(theta) = 1/sin(theta)

OpenStudy (anonymous):

wait um ok so how does it make sense? @math&ing001

OpenStudy (math&ing001):

Can you express csc(x) in terms of sin(x) ?

OpenStudy (anonymous):

idk? i dont even know how to do this?

OpenStudy (anonymous):

you said B made sense? thats the one we're talking about right?

OpenStudy (math&ing001):

Haven't you studied trigonometry yet ? We know that \(\tan(\Theta) = \frac{ \sin(\Theta) }{ \cos(\Theta) }\) and \(\cot(\Theta) = \frac{ 1 }{ \tan(\theta) }\) So \[\frac{ 1 }{ \tan(\Theta) }=\cos(\Theta) * (1/\sin(\Theta))=\cot(\Theta)\]

OpenStudy (anonymous):

I started a new school later in the year like a month ago and i wasnt on this at my old school so i like jumped so something completely different..

OpenStudy (math&ing001):

I guess you'll need extra classes for this, because you're gonna need this later.

OpenStudy (anonymous):

ya i'm taking more classes this summer to catch up